Find the next four terms of the arithmetic sequence 4,10,16,...
Thanks

Answer by venugopalramana(3286)

Find the next four terms of the arithmetic sequence 4,10,16,...
CD=10-4=16-10=6
HENCE NEXT 4 TERMS ARE
16+6=22
22+6=28
28+6=34
34+6=40